OpenStudy (anonymous):

(1+SINX)^(1/2) 1+sinx -------------= ------- (1-sinx)^(1/2) |cosx|

OpenStudy (anonymous):

verify

OpenStudy (anonymous):

okay, you have what exactly? this? \[\sqrt{\frac{1+sinx}{1-sin(x)}}=\frac{1+sin(x)}{|cos(x)|}\text{?}\]

OpenStudy (anonymous):

correct

OpenStudy (dumbcow):

multiply top and bottom by conjugate: sqrt(1+sin x) \[\frac{\sqrt{1+sinx}}{\sqrt{1-sinx}}*\frac{\sqrt{1+sinx}}{\sqrt{1+sinx}} = \frac{1+sinx}{\sqrt{1-\sin^{2} x}} = \frac{1+sinx}{cosx}\]

OpenStudy (campbell_st):

rationalise the denominator by multiplying by 1 + sin(x) will result in \[\sqrt{(1+\sin(x)^2/1 - \sin^2(x)}\] denominator is cos^2(x) while square root of the numerator is (1+sin(x)

OpenStudy (anonymous):

after the sqrt gets off the top wouldn't it go off the bottom as well

OpenStudy (campbell_st):

the reason its absolute value is that the square root will give a positive and negative answer.. hence the abs value